A Bibliography of Volumes in Lecture Notes in Computer Science (2011)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.02

Title word cross-reference

10th [AFG^{+}11, BS11a, BS11b, LPC^{+}11, LTW11, Pro11]. 12th [BC11, KK11b]. 13th [Che11, QSWL11]. 14th [KjHGG11, SM11, SBM11]. 15th [ALR11]. 16th [MK11]. 17th [LW11]. 18th [LZK11c, LZK11b, LZK11a]. 19th [CCMS11, RW11].

21st [KM11]. 22nd [AiNOW11, IS11]. 23rd [WZ11]. 24th [WR11].

37th [KK11a].

4th [AF11, JLS11a, JLS11b, Yan11a].

5th [Coe11, KM11, SAUL11, TGAH11, XW11, XL11].

6.3/7.3 [HHG11]. 6th [SLCT11].

75th [JL11]. 7th [CEK11, HS11, JM11, KMSB11, MRP11, SSO^{+}11, TKCW11b, TKCW11a].

8.8/11.2 [Pro11]. 8th [DKN^{+}11, HDBM11, SNRG11, Tho11].

9th [ABKT11, CTY^{+}11, KMMN11, Yan11b].

Aachen [JLS11a, JLS11b]. Abdominal [YC11]. ACM [KK11b]. ACM/ [KK11b]. Adaptive [DKN^{+}11]. ADG [SNRG11]. ADMA [TKCW11b, TKCW11a].

Advanced [Pro11, TKCW11b, TKCW11a]. Advances [SBM11, WR11, BS11a, BS11b, LW11, LPC^{+}11, TGAH11]. Agent [MRP11]. Agents [KjHGG11, KjHGG11]. AI [WR11]. AIRS [SSO^{+}11]. Aizu [KMSB11]. Aizu-Wakamatsu [KMSB11]. Algebraic [KR11]. Algorithms [IS11, AiNOW11]. Ambient [KMS^{+}11]. AMBN [OBM11]. AmI [KMS^{+}11]. AMR [DKN^{+}11]. Amsterdam [KMS^{+}11, MBH^{+}11, SL11]. Analysis [MK11, EVV^{+}11]. Anaphor [HDBM11]. Anaphora [HDBM11, HDBM11]. Andhra [PSDS11a, PSDS11b]. Answering [CTY^{+}11]. APLAS [Yan11b]. Application [LW11]. Applications [ABKT11, HDBM11, JLS11a, JLS11b, LC11, MOP11, MK11, Pro11, SM11, TKCW11b, TKCW11a]. Arab [SSO^{+}11]. Architectures [Ste11]. ArgMAS [MRP11]. Argumentation [MRP11]. Artificial [BS11a, BS11b, OBM11, SAUL11, WR11]. Asia [SSO^{+}11]. ASIACRYPT [LW11]. Asian [Yan11b]. Aspects [VEG^{+}11]. Athens [VEG^{+}11, XW11]. August [DKN^{+}11]. Australasian [WR11]. Australia [KjHGG11, WR11]. Austria [DKN^{+}11, HS11, HHG11]. Austrian [HS11]. Automated [SNRG11].

Based [LPC^{+}11, RW11, RW11]. BC [IS11]. Behavior [SL11]. Beijing [CY11, QSWL11, SLCT11, TKCW11b, TKCW11a, YC11]. Belgium [CTY^{+}11, Pro11]. Biometric [SLCT11]. Birthday [JL11]. Borrowdale [Tho11]. Bozapalidis [KR11]. Bratislava [KM11]. Brazil [SM11]. Brazilian [SM11]. Brian [JL11]. Budapest [EVV^{+}11].

CA [SBM11, XL11]. Cambridge [CCMS11]. Canada [AFG^{+}11, IS11, MRP11, STA^{+}11]. Canaria [TGAH11]. CANS [LTW11]. Card [Pro11]. CARDIS [Pro11]. Case [RW11, RW11]. Case-Based [RW11, RW11].

Castle [CMOZM11]. CCBR [SLCT11]. Celebration [HHG11]. Century [KM11]. Certified [JS11]. CGGA [ABKT11]. Challenges [HHG11, YC11]. Chennai [BC11]. Chile [MK11]. China [ABKT11, CY11, LPC^{+}11, LTW11, LC11, LZK11c, LZK11b, LZK11a, QSWL11, SLCT11, TKCW11b, TKCW11a, YC11]. Chinese [SLCT11]. CIARP [MK11]. Clinical [YC11]. Coding [Che11]. College [BKA11]. Colloquium [HDBM11]. Colonoscopy [YC11]. Combinatorial [IS11]. Communication [EVV^{+}11, HHG11]. Communications [QSWL11]. Compilers [Ste11]. Computation [AiNOW11]. Computational [CMOZM11, GTM11, YC11, ABKT11]. Computer [HS11, HHG11, KK11a, KR11, MK11, Tho11]. Computing [AFG^{+}11, BS11b, JL11, KMMN11, LC11, PSDS11a, PSDS11b]. Concepts [KK11a]. Confederated [MDH11]. Conference [ABKT11, AF11, AFG^{+}11, ALR11, BKA11, BS11a, BS11b, BC11, Che11, CY11, CTY^{+}11, Coe11, EVV^{+}11, HS11, JM11, JLS11a, JLS11b, JS11, KM11, KMMN11, KMS^{+}11, hKAS^{+}11, KjHGG11, KK11b, LW11, LPC^{+}11, LTW11, LC11, LZK11c, LZK11b, LZK11a, MOP11, MBH^{+}11, PSDS11a, PSDS11b, Pro11, QSWL11, RW11, SSO^{+}11, STA^{+}11, SLCT11, TKCW11b, TKCW11a, TGAH11, WR11, WZ11, XL11, ZRP^{+}11a, ZRP^{+}11b]. Congress [MK11]. Conjunction [hKAS^{+}11, YC11]. Context [DKN^{+}11]. Contributing [KM11]. COST [EVV^{+}11, VEG^{+}11]. CPP [JS11]. Crete [MDH11]. Critical [XW11]. CRITIS [XW11]. Cryptography [Yan11a, Che11]. Cryptology [BC11, LW11, LTW11]. Cyprus [KMMN11]. Czech [KK11a].

DAARC [HDBM11]. Dagstuhl [CMOZM11]. Dalian [ABKT11]. Data [SEK^{+}11, TKCW11b, TKCW11a]. Databases [KMSB11]. Datalog [dMGFS11, dMGFS11]. December [ALR11, AiNOW11, BS11a, BS11b, BC11, Che11, CEK11, CY11, JM11, JLS11a, JLS11b, JS11, KMMN11, KMSB11, hKAS^{+}11, KK11b, LW11, LPC^{+}11, LTW11, PSDS11a, PSDS11b, SSO^{+}11, STA^{+}11, SAUL11, SLCT11, TKCW11b, TKCW11a, WR11, XL11, Yan11b, Yan11a]. Decision [BKA11]. Dedicated [JL11, KR11]. Deduction [SNRG11]. Delaunay [GTM11]. Dependable [JL11]. Detection [SBM11]. Development [MOP11, RW11]. Diagrams [GTM11]. Digital [STA^{+}11]. disciplinary [SAUL11]. Discourse [HDBM11]. Distributed [ALR11]. DNIS [KMSB11]. Dubai [SSO^{+}11].

e-Health [HS11]. Economics [CEK11]. Edinburgh [AF11]. Education [KM11]. EI2N [MDH11]. Embedded [Ste11]. Emeritus [HHG11]. Emirates [SSO^{+}11]. Enactment [EVV^{+}11]. Engineering [HS11, Tho11, XL11]. Entertainment [AFG^{+}11]. EPEW [Tho11]. Essays [JL11, KR11]. European [Tho11]. Evaluation [HHG11]. Evolution [KM11]. Evolutionary [PSDS11a, PSDS11b]. Exploration [DKN^{+}11].

Faro [HDBM11]. FGIT [hKAS^{+}11]. First [JS11, dMGFS11]. Flexible [CTY^{+}11]. Formal [SM11]. Foundations [KR11, SM11]. Fourth [STA^{+}11]. FQAS [CTY^{+}11]. France [ALR11, WZ11]. Frontiers [OBM11]. Fusion [DKN^{+}11]. Future [HHG11, hKAS^{+}11].

Game [BKA11]. Games [AF11, MOP11]. GameSec [BKA11]. GDC [hKAS^{+}11]. Generation [hKAS^{+}11]. Geometry [ABKT11, SNRG11]. Germany [CMOZM11, JLS11a, JLS11b, SNRG11].

Ghent [CTY^{+}11]. Gran [TGAH11]. Graph [KK11a]. Graph-Theoretic [KK11a]. Graphs [ABKT11]. Graz [HS11]. Greece [MDH11, VEG^{+}11, XW11]. Günter [HHG11].

Haring [HHG11]. HBU [SL11]. Health [HS11]. Held [YC11]. Hersonissos [MDH11]. Hiding [SEK^{+}11]. High [Ste11]. High-Performance [Ste11]. Historic [JL11]. Hong [LPC^{+}11]. Honor [HHG11]. Human [HS11, SL11]. Human-Computer [HS11]. Hungary [EVV^{+}11]. Hyderabad [SAUL11].

Iberoamerican [MK11]. ICCBR [RW11]. ICE [MDH11]. ICEC [AFG^{+}11]. ICICA [LC11]. ICICS [QSWL11]. ICIDS [STA^{+}11]. ICIRA [JLS11a, JLS11b]. ICISS [JM11]. ICONIP [LZK11c, LZK11b, LZK11a]. ICSOC [KMMN11]. ICSP [MDH11]. ICSR [MBH^{+}11]. ICTSS [WZ11]. ICWL [LPC^{+}11]. IFIP [HHG11, Pro11, WZ11, KK11b]. IFIP/ [KK11b]. II [BS11b, JLS11b, LZK11b, PSDS11b, TKCW11b, ZRP^{+}11b]. III [LZK11c]. IMA [Che11]. IMACC [Che11]. Image [MK11]. Imaging [YC11]. INBAST [MDH11]. India [BC11, JM11, PSDS11a, PSDS11b, SAUL11]. INDOCRYPT [BC11]. Informatics [KM11, ZRP^{+}11a, ZRP^{+}11b, KM11]. Information [KMSB11, hKAS^{+}11, LW11, LZK11c, LZK11b, LZK11a, SSO^{+}11, XW11, HS11, JM11, LC11, QSWL11, SSO^{+}11]. Infrastructures [XW11]. Innovations [ZRP^{+}11a, ZRP^{+}11b]. Intelligence [BS11a, BS11b, KMS^{+}11, OBM11, SAUL11, WR11]. Intelligent [Coe11, JLS11a, JLS11b]. Interaction [HS11]. Interactive [STA^{+}11, STA^{+}11]. International [ABKT11, AF11, AFG^{+}11, ALR11, AiNOW11, BKA11, BS11a, BS11b, BC11, Che11, CEK11, CY11, CTY^{+}11, CCMS11, Coe11, CMOZM11, DKN^{+}11, EVV^{+}11, HHG11, IS11, JM11, JLS11a, JLS11b, JS11, KM11, KMMN11, KMS^{+}11, KMSB11, hKAS^{+}11, KjHGG11, KK11a, KK11b, LW11, LPC^{+}11, LTW11, LC11, LZK11c, LZK11b, LZK11a, MOP11, MRP11, MDH11, MBH^{+}11, PSDS11a, PSDS11b, Pro11, QSWL11, RW11, SL11, SNRG11, STA^{+}11, SAUL11, SBM11, TKCW11b, TKCW11a, TGAH11, VEG^{+}11, WZ11, XW11, XL11, Yan11a, YC11, ZRP^{+}11a, ZRP^{+}11b, dMGFS11]. Internet [MDH11, CEK11]. Intrusion [SBM11]. INTRUST [CY11]. Invited [MRP11]. Irvine [XL11]. ISAAC [AiNOW11]. isAI [OBM11]. ISDE [MDH11]. Island [hKAS^{+}11]. ISS [OBM11]. ISSEP [KM11]. Issue [GTM11]. Issues [EVV^{+}11]. Italy [Coe11]. IV [Ste11]. IVIC [ZRP^{+}11a, ZRP^{+}11b]. IWOCA [IS11].

January [Coe11]. Japan [AiNOW11, KMSB11, OBM11]. Jeju [hKAS^{+}11]. Joint [KMS^{+}11, WR11]. JSAI [OBM11]. JSAI-isAI [OBM11]. July [IS11, SNRG11]. June [KK11a]. JURISIN [OBM11].

Kenting [JS11, Yan11b]. Knowledge [XL11]. Kolkata [JM11]. Kong [LPC^{+}11]. Korea [hKAS^{+}11, LW11]. KSEM [XL11].

Languages [Yan11b]. Learning [LPC^{+}11, Coe11]. LENLS [OBM11]. Leuven [Pro11]. Linz [DKN^{+}11]. LION [Coe11]. Lisbon [KK11b, MOP11]. London [RW11].

Malaysia [ZRP^{+}11a, ZRP^{+}11b]. Management [XL11]. March [CCMS11, dMGFS11]. Maryland [BKA11]. May [MRP11]. MD [BKA11]. Meaningful [MDH11]. Memetic [PSDS11a, PSDS11b]. Menlo [SBM11]. Methods [SM11].

Mexican [BS11a, BS11b]. Mexico [BS11a, BS11b]. MICAI [BS11a, BS11b]. MICCAI [YC11]. Middleware [KK11b, KK11b]. MIG [AF11]. Milestones [HHG11]. Mining [TKCW11b, TKCW11a]. MIWAI [SAUL11]. Monastery [KK11a]. MONET [MDH11]. Motion [AF11]. Move [MDH11]. Multi [MRP11, SAUL11]. Multi-Agent [MRP11]. Multi-disciplinary [SAUL11]. Multidisciplinary [VEG^{+}11]. Multimedia [DKN^{+}11, SEK^{+}11]. Munich [SNRG11].

Netherlands [KMS^{+}11, MBH^{+}11, SL11]. Network [CEK11, LTW11]. Networked [KMSB11]. Neural [LZK11c, LZK11b, LZK11a]. NOLISP [TGAH11]. Nonlinear [TGAH11]. Nonverbal [EVV^{+}11]. November [ABKT11, AF11, BKA11, BS11a, BS11b, HS11, KMS^{+}11, KjHGG11, LZK11c, LZK11b, LZK11a, MK11, MBH^{+}11, OBM11, QSWL11, SL11, STA^{+}11, TGAH11, WZ11, Yan11a, ZRP^{+}11a, ZRP^{+}11b]. NSF [MDH11].

Occasion [HHG11, JL11, KR11]. October [AFG^{+}11, CTY^{+}11, CMOZM11, HDBM11, HHG11, KM11, LC11, MDH11, Tho11, VEG^{+}11]. OPODIS [ALR11]. Opportunities [YC11]. Optimization [Coe11]. Oriented [KMMN11]. ORM [MDH11]. OTM [MDH11]. OTMA [MDH11]. Oxford [Che11, dMGFS11].

Palmas [TGAH11]. Papers [ABKT11, CY11, CCMS11, Coe11, CMOZM11, DKN^{+}11, EVV^{+}11, HDBM11, HHG11, IS11, KK11a, MRP11, OBM11, Pro11, SNRG11, SM11, VEG^{+}11, XW11, YC11, dMGFS11]. Paphos [KMMN11]. Paris [WZ11]. Park [BKA11, SBM11]. Part [BS11a, BS11b, JLS11a, JLS11b, LZK11c, LZK11b, LZK11a, PSDS11a, PSDS11b, TKCW11b, TKCW11a, ZRP^{+}11a, ZRP^{+}11b]. Pattern [MK11]. Paulo [SM11]. Perception [VEG^{+}11]. PERFORM [HHG11]. Performance [Ste11, Tho11, HHG11]. Perspectives [KM11]. Perth [WR11]. Portugal [HDBM11, KK11b, MOP11]. Post [Yan11a]. Post-Quantum [Yan11a]. Posters [MDH11]. PQCrypto [Yan11a]. Practice [KjHGG11]. Pradesh [PSDS11a, PSDS11b]. PRIMA [KjHGG11]. Principle [KjHGG11]. Principles [ALR11]. Procedings [JM11]. Proceedings [BS11a, BS11b, JLS11a, JLS11b, LZK11c, LZK11b, LZK11a, PSDS11a, PSDS11b, TKCW11b, TKCW11a, ZRP^{+}11a, ZRP^{+}11b, AF11, AFG^{+}11, ALR11, AiNOW11, BKA11, BC11, Che11, CEK11, CTY^{+}11, HS11, JS11, KM11, KMMN11, KMS^{+}11, KMSB11, hKAS^{+}11, KjHGG11, KK11b, LW11, LPC^{+}11, LTW11, LC11, MOP11, MK11, MDH11, MBH^{+}11, QSWL11, RW11, SL11, SSO^{+}11, STA^{+}11, SAUL11, SBM11, SLCT11, Tho11, TGAH11, WR11, WZ11, XL11, Yan11b, Yan11a]. Processing [CMOZM11, EVV^{+}11, HDBM11, LZK11c, LZK11b, LZK11a, TGAH11]. Programming [Yan11b]. Programs [JS11]. Progress [BC11, MK11]. Proofs [JS11]. Protocols [CCMS11]. Pucón [MK11]. Puebla [BS11a, BS11b].

Qinhuangdao [LC11]. Quality [HS11]. Quantum [Yan11a]. Query [CTY^{+}11].

RAID [SBM11]. Randell [JL11]. Reasoning [RW11]. Recognition [MK11, SLCT11]. Reloaded [dMGFS11]. Republic [KK11a]. Research [Pro11, RW11, ZRP^{+}11a, ZRP^{+}11b]. Resolution [HDBM11]. Retirement [KR11]. Retrieval [DKN^{+}11, SSO^{+}11]. Revised [ABKT11, CY11, CCMS11, CMOZM11, DKN^{+}11, EVV^{+}11, HDBM11, HHG11, IS11, KK11a, MRP11, OBM11, Pro11, SNRG11, SM11, VEG^{+}11, XW11, YC11, dMGFS11]. Robotics [JLS11a, JLS11b, MBH^{+}11]. Rome [Coe11].

Sanya [LTW11]. São [SM11]. SBMF [SM11]. Schools [KM11]. Science [GTM11, KK11a, KR11, XL11]. Second [BKA11, CY11, KMS^{+}11, LC11, MOP11, PSDS11a, PSDS11b, SL11, YC11, ZRP^{+}11a, ZRP^{+}11b]. Security [BKA11, JM11, LW11, LTW11, QSWL11, SEK^{+}11, XW11, CCMS11]. SeDeS [MDH11]. Selangor [ZRP^{+}11a, ZRP^{+}11b]. Selected [ABKT11, CY11, CCMS11, Coe11, DKN^{+}11, EVV^{+}11, HDBM11, HHG11, IS11, MRP11, OBM11, Pro11, SNRG11, SM11, VEG^{+}11, YC11, dMGFS11]. SEMCCO [PSDS11a, PSDS11b]. Seminar [CMOZM11]. Seoul [LW11]. September [EVV^{+}11, MOP11, Pro11, RW11, SM11, SBM11, XW11, YC11]. Serious [MOP11]. Service [KMMN11]. Service-Oriented [KMMN11]. SGDA [MOP11]. Shanghai [LZK11c, LZK11b, LZK11a]. Singapore [CEK11]. Situation [KM11]. Slovakia [KM11]. Smart [Pro11]. Social [MBH^{+}11]. Societies [SSO^{+}11]. Society [HS11]. Soft [BS11b]. Software [WZ11]. South [LW11]. Spain [TGAH11]. Special [GTM11]. Speech [TGAH11]. Storytelling [STA^{+}11]. Sustaining [ZRP^{+}11a, ZRP^{+}11b]. Swarm [PSDS11a, PSDS11b]. SWWS [MDH11]. Symeon [KR11]. Symposium [AiNOW11, SM11, SBM11, Yan11b]. Systems [ALR11, CY11, CTY^{+}11, HHG11, JM11, KMSB11, MRP11, MDH11, WZ11, Yan11b].

Taipei [Yan11a]. Taiwan

[JS11, Yan11b, Yan11a]. TD0904 [VEG^{+}11]. Technology [hKAS^{+}11, SSO^{+}11]. Teplá [KK11a]. Testing [WZ11]. Theoretic [KK11a]. Theory [BKA11, LW11]. Third [hKAS^{+}11, MBH^{+}11]. Time [VEG^{+}11]. Tokyo [OBM11]. Toronto [MRP11]. Toulouse [ALR11]. Transactions [GTM11, SEK^{+}11, Ste11]. Trends [SAUL11]. Triangulation [GTM11]. Trusted [CY11].

UK [AF11, Che11, CCMS11, RW11, Tho11, dMGFS11]. Understanding [SL11]. United [SSO^{+}11]. USA [BKA11, SBM11, XL11]. USAB [HS11]. Usability [HS11]. USENIX [KK11b].

VADER [MDH11]. Vancouver [AFG^{+}11, STA^{+}11]. Verbal [EVV^{+}11]. VI [SEK^{+}11]. Victoria [IS11]. Video [CMOZM11, CMOZM11]. Vienna [HHG11]. Virtual [YC11]. Visakhapatnam [PSDS11a, PSDS11b]. Vision [MK11]. Visual [ZRP^{+}11a, ZRP^{+}11b, ZRP^{+}11a, ZRP^{+}11b]. Voronoi [GTM11].

Wakamatsu [KMSB11]. Web [LPC^{+}11]. Web-Based [LPC^{+}11]. WG [HHG11, KK11a, Pro11, WZ11]. WINE [CEK11]. Wollongong [KjHGG11]. Workgroup [HS11]. Workshop [CEK11, CCMS11, DKN^{+}11, HHG11, IS11, KMSB11, KK11a, MRP11, SL11, SNRG11, SAUL11, Tho11, VEG^{+}11, XW11, Yan11a, YC11, dMGFS11]. Workshops [MDH11, OBM11].

XIV [GTM11]. XIX [CCMS11].

Yokohama [AiNOW11].

References

Akiyama:2011:CGG

[ABKT11] Jin Akiyama, Jiang Bo, Mikio Kano, and Xuehou Tan, editors. Computational Geometry, Graphs and Applications: 9th International Conference, CGGA 2010, Dalian, China, November 3–6, Revised Selected Papers, volume 7033 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24982-5. ISSN 0302-9743 (print), 1611-3349 (elec-

                                                   tronic).  URL http://www.
                                                   springerlink.com/content/
                                                   978-3-642-24982-2.

Allbeck:2011:MGI

[AF11] Jan M. Allbeck and Petros Faloutsos, editors. Motion in Games: 4th International Conference, MIG 2011, Edinburgh, UK, November 13– 15. Proceedings, volume 7060 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25089-0. ISSN 0302-9743 (print), 1611-3349 (elec-

                                                   tronic).  URL http://www.
                                                   springerlink.com/content/
                                                   978-3-642-25089-7.

Anacleto:2011:ECI

[AFG^{+}11] Junia Coutinho Anacleto, Sidney Fels, Nicholas Graham, Bill Kapralos, Magy Saif El-Nasr, and et al, editors. Entertainment Computing — ICEC 2011: 10th International Conference, ICEC 2011, Vancouver, Canada, October 5–8. Proceedings, volume 6972 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-24499-8. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-24499-5.

Asano:2011:ACI

[AiNOW11] Takao Asano, Shin ichi Nakano, Yoshio Okamoto, and Osamu Watanabe, editors. Algorithms and Computation: 22nd International Symposium, ISAAC 2011, Yokohama, Japan, December 5–8. Proceedings, volume 7074 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25590-6. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25590-8.

Anta:2011:PDS

[ALR11] Antonio Fernàndez Anta, Giuseppe Lipari, and Matthieu Roy, editors. Principles of Distributed Systems: 15th International Conference, OPODIS 2011, Toulouse, France, December 13–16. Proceedings, volume 7109 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9.

ISBN 3-642-25872-7. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-25872-5.

Bernstein:2011:PCI

[BC11] Daniel J. Bernstein and Sanjit Chatterjee, editors. Progress in Cryptology — INDOCRYPT 2011: 12th International Conference on Cryptology in India, Chennai, India, December 11–14. Proceedings, volume 7107 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25577-9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-25577-9.

Baras:2011:DGT

[BKA11] John S. Baras, Jonathan Katz, and Eitan Altman, editors. Decision and Game Theory for Security: Second International Conference, GameSec 2011, College Park, MD, Maryland, USA, November 14–15. Proceedings, volume 7037 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-25279-6. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-25279-2.

Batyrshin:2011:AAI

[BS11a] Ildar Batyrshin and Grigori Sidorov, editors. Advances in Artificial Intelligence: 10th Mexican International Conference on Artificial Intelligence, MICAI 2011, Puebla, Mexico, November 26 — December 4, 2011, Proceedings, Part I, volume 7094 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25323-7. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25323-2.

Batyrshin:2011:ASC

[BS11b] Ildar Batyrshin and Grigori Sidorov, editors. Advances in Soft Computing: 10th Mexican International Conference on Artificial Intelligence, MICAI 2011, Puebla, Mexico, November 26 — December 4, 2011, Proceedings, Part II, volume 7095 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25329-6. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25329-4.

Christianson:2011:SPX

[CCMS11] Bruce Christianson, Bruno Crispo, James Malcolm, and Frank Stajano, editors. Security Protocols XIX: 19th International Workshop, Cambridge, UK, March 28–30, Revised Selected Papers, volume 7114 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25866-2. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25866-4.

Chen:2011:INE

[CEK11] Ning Chen, Edith Elkind, and Elias Koutsoupias, editors. Internet and Network Economics: 7th International Workshop, WINE 2011, Singapore, December 11–14. Proceedings, volume 7090 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25509-4. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25509-0.

Chen:2011:CCI

[Che11] Liqun Chen, editor. Cryptography and Coding: 13th IMA International Conference, IMACC 2011, Oxford, UK, December 12–15. Proceedings, volume 7089 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25515-9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.

springerlink.com/content/
978-3-642-25515-1.

Cremers:2011:VPC

[CMOZM11] Daniel Cremers, Marcus Magnor, Martin R. Oswald, and Lihi Zelnik-Manor, editors. Video Processing and Computational Video: International Seminar, Dagstuhl Castle, Germany, October 10–15. Revised Papers, volume 7082 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24869-1. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-24869-6.

Coello:2011:LIO

[Coe11] Carlos A. Coello Coello, editor. Learning and Intelligent Optimization: 5th International Conference, LION 5, Rome, Italy, January 17– 21. Selected Papers, volume 6683 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25565-5. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25565-6.

Christiansen:2011:FQA

[CTY^{+}11] Henning Christiansen, Guy De Tré, Adnan Yazici, Slawomir Zadrozny, Troels Andreasen, and et al, editors. Flexible Query Answering Systems: 9th International Conference, FQAS 2011, Ghent, Belgium, October 26–28 Proceedings, volume 7022 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24763-6. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-24763-7.

Chen:2011:TSS

[CY11] Liqun Chen and Moti Yung, editors. Trusted Systems: Second International Conference, INTRUST 2010, Beijing, China, December 13–15, Revised Selected Papers, volume 6802 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25282-6. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25282-2.

Detyniecki:2011:AMR

[DKN^{+}11] Marcin Detyniecki, Peter Knees, Andreas Nürnberger, Markus Schedl, and Sebastian Stober, editors. Adaptive Multimedia Retrieval. Context, Exploration, and Fusion: 8th International Workshop, AMR 2010, Linz, Austria, August 17–18, Revised Selected Papers, volume 6817 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9.

ISBN 3-642-27168-5. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-27168-7.

deMoor:2011:DRF

[dMGFS11] Oege de Moor, Georg Gottlob, Tim Furche, and Andrew Sellers, editors. Datalog Reloaded: First International Workshop, Datalog 2010, Oxford, UK, March 16–19. Revised Selected Papers, volume 6702 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24205-7. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-24205-2.

Esposito:2011:AVN

[EVV^{+}11] Anna Esposito, Alessandro Vinciarelli, Klára Vicsi, Catherine Pelachaud, and Anton Nijholt, editors. Analysis of Verbal and Nonverbal Communication and Enactment. The Processing Issues: COST 2102 International Conference, Budapest, Hungary, September 7– 10, Revised Selected Papers, volume 6800 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25774-7. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25774-2.

Gavrilova:2011:TCS

[GTM11] Marina L. Gavrilova, C. J. Kenneth Tan, and Mir Abolfazl Mostafavi, editors. Transactions on Computational Science XIV: Special Issue on Voronoi Diagrams and Delaunay Triangulation, volume 6970 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25248-6. ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        ???? pp. URL http://www.
                                        springerlink.com/content/
                                        978-3-642-25248-8.

Hendrickx:2011:APA

[HDBM11] Iris Hendrickx, Sobha Lalitha Devi, António Branco, and Ruslan Mitkov, editors. Anaphora Processing and Applications: 8th Discourse Anaphora and Anaphor Resolution Colloquium, DAARC 2011, Faro, Portugal, October 6–7. Revised Selected Papers, volume 7099 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25916-2. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-25916-6.

Hummel:2011:PEC

[HHG11] Karin Anna Hummel, Helmut Hlavacs, and Wilfried Gansterer, editors. Performance Evaluation of Computer and Communication Systems. Milestones and Future Challenges: IFIP WG 6.3/7.3 International Workshop, PERFORM 2010, in Honor of Günter Haring on the Occasion of His Emeritus Celebration, Vienna, Austria, October 14–16, Revised Selected Papers, volume 6821 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25574-4. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25574-8.

Kim:2011:FGI

[hKAS^{+}11] Tai hoon Kim, Hojjat Adeli, Dominik Slezak, Frode Eika Sandnes, Xiaofeng Song, and et al, editors. Future Generation Information Technology: Third International Conference, FGIT 2011 in Conjunction with GDC 2011, Jeju Island, Korea, December 8–10. Proceedings, volume 7105 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-27141-3. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-27141-0.

Holzinger:2011:IQH

[HS11] Andreas Holzinger and Klaus-Martin Simonic, editors. Information Quality in e-Health: 7th Conference of the Workgroup Human-Computer Interaction and Usability Engineering of the Austrian Computer Society, USAB 2011, Graz, Austria, November 25–26. Proceedings, volume 7058 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25363-6. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-25363-8.

Iliopoulos:2011:CAI

[IS11] Costas S. Iliopoulos and William F. Smyth, editors. Combinatorial Algorithms: 22nd International Workshop, IWOCA 2011, Victoria, BC, Canada, July 20– 22, Revised Selected Papers, volume 7056 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25010-6. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-25010-1.

Jones:2011:DHC

[JL11] Cliff B. Jones and John L. Lloyd, editors. Dependable and Historic Computing: Essays Dedicated to Brian Randell on the Occasion of His 75th Birthday, volume 6875 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24540-4, 3-642-24541-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

???? pp. URL http://www.
springerlink.com/content/
978-3-642-24540-4.

Jeschke:2011:IRAa

[JLS11a] Sabina Jeschke, Honghai Liu, and Daniel Schilberg, editors. Intelligent Robotics and Applications: 4th International Conference, ICIRA 2011, Aachen, Germany, December 6–8, Proceedings, Part I, volume 7101 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25485-3. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25485-7.

Jeschke:2011:IRAb

[JLS11b] Sabina Jeschke, Honghai Liu, and Daniel Schilberg, editors. Intelligent Robotics and Applications: 4th International Conference, ICIRA 2011, Aachen, Germany, December 6–8, Proceedings, Part II, volume 7102 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25488-8. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25488-8.

Jajodia:2011:ISS

[JM11] Sushil Jajodia and Chandan Mazumdar, editors. Information Systems Security: 7th International Conference, ICISS 2011, Kolkata, India, December 15–19, Procedings, volume 7093 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25559-0. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25559-5.

Jouannaud:2011:CPP

[JS11] Jean-Pierre Jouannaud and Zhong Shao, editors. Certified Programs and Proofs: First International Conference, CPP 2011, Kenting, Taiwan, December 7–9. Proceedings, volume 7086 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25378-4. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25378-2.

Kinny:2011:APA

[KjHGG11] David Kinny, Jane Yung jen Hsu, Guido Governatori, and Aditya K. Ghose, editors. Agents in Principle, Agents in Practice: 14th International Conference, PRIMA 2011, Wollongong, Australia, November 16–18. Proceedings, volume 7047 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25043-2. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25043-9.

Kolman:2011:GTC

[KK11a] Petr Kolman and Jan Kratochvíl, editors. Graph-Theoretic Concepts in Computer Science: 37th International Workshop, WG 2011, Teplá Monastery, Czech Republic, June 21–24. Revised Papers, volume 6986 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25869-7. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25869-5.

Kon:2011:MAI

[KK11b] Fabio Kon and Anne-Marie Kermarrec, editors. Middleware 2011: ACM/ IFIP/ USENIX 12th International Middleware Conference, Lisbon, Portugal, December 12–16. Proceedings, volume 7049 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25820-4. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25820-6.

Kalas:2011:ISC

[KM11] Ivan Kalas and Roland T. Mittermeir, editors. Informatics in Schools. Contributing to 21st Century Education: 5th International Conference on Informatics in Schools: Situation, Evolution and Perspectives, IS-SEP 2011, Bratislava, Slovakia, October 26–29. Proceedings, volume 7013 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24721-0. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-24721-7.

Kappel:2011:SOC

[KMMN11] Gerti Kappel, Zakaria Maamar, and Hamid R. Motahari-Nezhad, editors. Service-Oriented Computing: 9th International Conference, ICSOC 2011, Paphos, Cyprus, December 5–8 Proceedings, volume 7084 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25534-5. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25534-2.

Keyson:2011:AIS

[KMS^{+}11] David V. Keyson, Mary Lou Maher, Norbert Streitz, Adrian Cheok, Juan Carlos Augusto, and et al, editors. Ambient Intelligence: Second International Joint Conference on AmI 2011, Amsterdam, The Netherlands, November 16–18. Proceedings, volume 7040 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25166-8. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25166-5.

Kikuchi:2011:DNI

[KMSB11] Shinji Kikuchi, Aastha Madaan, Shelly Sachdeva, and Subhash Bhalla, editors. Databases in Networked Information Systems: 7th International Workshop, DNIS 2011, Aizu-Wakamatsu, Japan, December 12–14. Proceedings, volume 7108 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-25730-5. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25730-8.

Kuich:2011:AFC

[KR11] Werner Kuich and George Rahonis, editors. Algebraic Foundations in Computer Science: Essays Dedicated to Symeon Bozapalidis on the Occasion of His Retirement, volume 7020 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24896-9. ISSN 0302-9743 (print), 1611-3349 (electronic).

???? pp. URL http://www.

                                        springerlink.com/content/
                                        978-3-642-24896-2.

Liu:2011:ICA

[LC11] Baoxiang Liu and Chunlai Chai, editors. Information Computing and Applications: Second International Conference, ICICA 2011, Qinhuangdao, China, October 28–31. Proceedings, volume 7030 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25254-0. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-25254-9.

Leung:2011:AWB

[LPC^{+}11] Howard Leung, Elvira Popescu, Yiwei Cao, Rynson W. H. Lau, and Wolfgang Nejdl, editors. Advances in Web-Based Learning — ICWL 2011: 10th International Conference, Hong Kong, China, December 8– 10. Proceedings, volume 7048 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25812-3. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-25812-1.

Lin:2011:CNS

[LTW11] Dongdai Lin, Gene Tsudik, and Xiaoyun Wang, editors. Cryptology and Network Security: 10th International Conference, CANS 2011, Sanya, China, December 10–12. Proceedings, volume 7092 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25512-4. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25512-0.

Lee:2011:ACA

[LW11] Dong Hoon Lee and Xiaoyun Wang, editors. Advances in Cryptology — ASI-ACRYPT 2011: 17th International Conference on the Theory and Application of Cryptology and Information Security, Seoul, South Korea, December 4–8. Proceedings, volume 7073 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25384-9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25384-3.

Lu:2011:NIPc

[LZK11a] Bao-Liang Lu, Liqing Zhang, and James Kwok, editors. Neural Information Processing: 18th International Conference, ICONIP 2011, Shanghai, China, November 13–17, Proceedings, Part I, volume 7062 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24954-X. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-24954-9.

Lu:2011:NIPb

[LZK11b] Bao-Liang Lu, Liqing Zhang, and James Kwok, editors. Neural Information Processing: 18th International Conference, ICONIP 2011, Shanghai, China, November 13–17, Proceedings, Part II, volume 7063 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24957-4. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-24957-0.

Lu:2011:NIPa

[LZK11c] Bao-Liang Lu, Liqing Zhang, and James Kwok, editors. Neural Information Processing: 18th International Conference, ICONIP 2011, Shanghai, China, November 13–17, Proceedings, Part III, volume 7064 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24964-7. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-24964-8.

Mutlu:2011:SRT

[MBH^{+}11] Bilge Mutlu, Christoph Bartneck, Jaap Ham, Vanessa Evers, and Takayuki Kanda, editors. Social Robotics: Third International Conference, ICSR 2011, Amsterdam, The Netherlands, November 24–25. Proceedings, volume 7072 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25503-5. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25503-8.

Meersman:2011:MMI

[MDH11] Robert Meersman, Tharam Dillon, and Pilar Herrero, editors. On the Move to Meaningful Internet Systems: OTM 2011 Workshops: Confederated International Workshops and Posters: EI2N+NSF ICE, ICSP+INBAST, ISDE, ORM, OTMA, SWWS+MONET+SeDeS, and VADER 2011, Hersonissos, Crete, Greece, October 17– 21. Proceedings, volume 7046 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25125-0. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25125-2.

Martin:2011:PPR

[MK11] César San Martin and Sang-Woon Kim, editors. Progress in Pattern Recognition, Image Analysis, Computer Vision, and Applications: 16th Iberoamerican Congress, CIARP 2011, Pucón, Chile, November 15– 18. Proceedings, volume 7042 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25084-X. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-25084-2.

Ma:2011:SGD

[MOP11] Minhua Ma, Manuel Fradinho Oliveira, and João Madeiras Pereira, editors. Serious Games Development and Applications: Second International Conference, SGDA 2011, Lisbon, Portugal, September 19–20. Proceedings, volume 6944 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-23833-5. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-23833-8.

McBurney:2011:AMA

[MRP11] Peter McBurney, Iyad Rahwan, and Simon Parsons, editors. Argumentation in Multi-Agent Systems: 7th International Workshop, ArgMAS 2010 Toronto, ON, Canada, May 10, 2010 Revised, Selected and Invited Papers, volume 6614 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21939-X. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-21939-9.

Onada:2011:NFA

[OBM11] Takashi Onada, Daisuke Bekki, and Eric McCready, editors. New Frontiers in Artificial Intelligence: JSAI-isAI 2010 Workshops, LENLS, JURISIN, AMBN, ISS, Tokyo, Japan, November 18–19, Revised Selected Papers, volume 6797 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25654-6. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25654-7.

Prouff:2011:SCR

[Pro11] Emmanuel Prouff, editor. Smart Card Research and Advanced Applications: 10th IFIP WG 8.8/11.2 International Conference, CARDIS 2011, Leuven, Belgium, September 14–16, Revised Selected Papers, volume 7079 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-27256-8. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-27256-1.

Panigrahi:2011:SEMa

[PSDS11a] Bijaya Ketan Panigrahi, Ponnuthurai Nagaratnam Suganthan, Swagatam Das, and Suresh Chandra Satapathy, editors. Swarm, Evolutionary, and Memetic Computing: Second International Conference, SEM-CCO 2011, Visakhapatnam, Andhra Pradesh, India, December 19–21, Proceedings, Part I, volume 7076 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-27171-5. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-27171-7.

Panigrahi:2011:SEMb

[PSDS11b] Bijaya Ketan Panigrahi, Ponnuthurai Nagaratnam Suganthan, Swagatam Das, and Suresh Chandra Satapathy, editors. Swarm, Evolutionary, and Memetic Computing: Second International Conference, SEM-CCO 2011, Visakhapatnam, Andhra Pradesh, India, December 19–21, Proceedings, Part II, volume 7077 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-27241-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

                                        springerlink.com/content/
                                        978-3-642-27241-7.

Qing:2011:ICS

[QSWL11] Sihan Qing, Willy Susilo, Guilin Wang, and Dongmei Liu, editors. Information and Communications Security: 13th International Conference, ICICS 2011, Beijing, China, November 23–26. Proceedings, volume 7043 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25242-7. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25242-6.

Ram:2011:CBR

[RW11] Ashwin Ram and Nirmalie Wiratunga, editors. Case-Based Reasoning Research and Development: 19th International Conference on Case-Based Reasoning, ICCBR 2011, London, UK, September 12– 15. Proceedings, volume 6880 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23290-6. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-23290-9.

Sombattheera:2011:MDT

[SAUL11] Chattrakul Sombattheera, Arun Agarwal, Siba K. Udgata, and Kittichai Lavangnananda, editors. Multi-disciplinary Trends in Artificial Intelligence: 5th International Workshop, MIWAI 2011, Hyderabad, India, December 7–9. Proceedings, volume 7080 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25724-0. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25724-7.

Sommer:2011:RAI

[SBM11] Robin Sommer, Davide Balzarotti, and Gregor Maier, editors. Recent Advances in Intrusion Detection: 14th International Symposium, RAID 2011, Menlo Park, CA, USA, September 20– 21. Proceedings, volume 6961 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-23643-X. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-23643-3.

Shi:2011:TDH

[SEK^{+}11] Yun Q. Shi, Sabu Emmanuel, Mohan S. Kankanhalli, Shih-Fu Chang, Regunathan Radhakrishnan, and et al, editors. Transactions on Data Hiding and Multimedia Security VI, volume 6730 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24555-2. ISSN 0302-9743 (print), 1611-3349 (electronic).

???? pp. URL http://www.
springerlink.com/content/
978-3-642-24555-8.

Salah:2011:HBU

[SL11] Albert Ali Salah and Bruno Lepri, editors. Human Behavior Understanding: Second International Workshop, HBU 2011, Amsterdam, The Netherlands, November 16, 2011. Proceedings, volume 7065 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-25445-4. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25445-1.

Sun:2011:BRC

[SLCT11] Zhenan Sun, Jianhuang Lai, Xilin Chen, and Tieniu Tan, editors. Biometric Recognition: 6th Chinese Conference, CCBR 2011, Beijing, China, December 3–4. Proceedings, volume 7098 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25448-9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25448-2.

Simao:2011:FMF

[SM11] Adenilso Simao and Carroll Morgan, editors. Formal Methods, Foundations and Applications: 14th Brazilian Symposium, SBMF 2011, São Paulo, Brazil, September 26–30, Revised Selected Papers, volume 7021 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25031-9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25031-6.

Schreck:2011:ADG

[SNRG11] Pascal Schreck, Julien Narboux, and Jürgen Richter-Gebert, editors. Automated Deduction in Geometry: 8th International Workshop, ADG 2010, Munich, Germany, July 22–24, Revised Selected Papers, volume 6877 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25069-6. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25069-9.

Salem:2011:IRT

[SSO^{+}11] Mohamed VallMohamed Salem, Khaled Shaalan, Farhad Oroumchian, Azadeh Shakery, and Halim Khelalfa, editors. Information Retrieval Technology: 7th Asia Information Retrieval Societies Conference, AIRS 2011, Dubai, United Arab Emirates, December 18– 20. Proceedings, volume 7097 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25630-9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25630-1.

Si:2011:ISF

[STA^{+}11] Mei Si, David Thue, Elisabeth André, James C. Lester, Joshua Tanenbaum, and et al, editors. Interactive Storytelling: Fourth International Conference on Interactive Digital Storytelling, ICIDS 2011, Vancouver, Canada, November 28 — 1 December, 2011. Proceedings, volume 7069 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25288-5. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25288-4.

Stenstrom:2011:THP

[Ste11] Per Stenström, editor. Transactions on High-Performance Embedded Architectures and Compilers IV, volume 6760 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24567-6. ISSN 0302-9743 (print), 1611-3349 (electronic).

???? pp. URL http://www.
springerlink.com/content/
978-3-642-24567-1.

Travieso-Gonzalez:2011:ANS

[TGAH11] Carlos M. Travieso-González and Jesús B. Alonso-Hernández, editors. Advances in Nonlinear Speech Processing: 5th International Conference on Nonlinear Speech Processing, NO-LISP 2011, Las Palmas de Gran Canaria, Spain, November 7–9. Proceedings, volume 7015 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25019-X. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-25019-4.

Thomas:2011:CPE

[Tho11] Nigel Thomas, editor. Computer Performance Engineering: 8th European Performance Engineering Workshop, EPEW 2011, Borrowdale, UK, October 12–13. Proceedings, volume 6977 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24748-2. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-24748-4.

Tang:2011:ADMb

[TKCW11a] Jie Tang, Irwin King, Ling Chen, and Jianyong Wang, editors. Advanced Data Mining and Applications: 7th International Conference, ADMA 2011, Beijing, China, December 17–19, Proceedings, Part I, volume 7120 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25852-2. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25852-7.

Tang:2011:ADMa

[TKCW11b] Jie Tang, Irwin King, Ling Chen, and Jianyong Wang, editors. Advanced Data Mining and Applications: 7th International Conference, ADMA 2011, Beijing, China, December 17–19, Proceedings, Part II, volume 7121 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25855-7. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25855-8.

Vatakis:2011:MAT

[VEG^{+}11] Argiro Vatakis, Anna Esposito, Maria Giagkou, Fred Cummins, and Georgios Papadelis, editors. Multidisciplinary Aspects of Time and Time Perception: COST TD0904 International Workshop, Athens, Greece, October 7–8, Revised Selected Papers, volume 6789 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21477-0. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-21477-6.

Wang:2011:AAA

[WR11] Dianhui Wang and Mark Reynolds, editors. AI 2011: Advances in Artificial Intelligence: 24th Australasian Joint Conference, Perth, Australia, December 5–8. Proceedings, volume 7106 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25831-X. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-25831-2.

Wolff:2011:TSS

[WZ11] Burkhart Wolff and Fatiha Zaïdi, editors. Testing Software and Systems: 23rd IFIP WG 6.1 International Conference, ICTSS 2011, Paris, France, November 7–10. Proceedings, volume 7019 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-24579-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

                                        springerlink.com/content/
                                        978-3-642-24579-4.

Xiong:2011:KSE

[XL11] Hui Xiong and W. B. Lee, editors. Knowledge Science, Engineering and Management:

5th International Conference, KSEM 2011, Irvine, CA, USA, December 12–14. Proceedings, volume 7091 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25974-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

springerlink.com/content/
978-3-642-25974-6.

Xenakis:2011:CII

[XW11] Christos Xenakis and Stephen Wolthusen, editors. Critical Information Infrastructures Security: 5th International Workshop, CRITIS 2010, Athens, Greece, September 23–24. Revised Papers, volume 6712 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-21693-5. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-21693-0.

Yang:2011:PQC

[Yan11a] Bo-Yin Yang, editor. Post-Quantum Cryptography: 4th International Workshop, PQCrypto 2011, Taipei, Taiwan, November 29 — December 2, 2011. Proceedings, volume 7071 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25404-7. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-25404-8.

Yang:2011:PLS

[Yan11b] Hongseok Yang, editor. Programming Languages and Systems: 9th Asian Symposium, APLAS 2011, Kenting, Taiwan, December 5–7. Proceedings, volume 7078 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25317-2. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-25317-1.

Yoshida:2011:VCA

[YC11] Hiroyuki Yoshida and Wenli Cai, editors. Virtual Colonoscopy and Abdominal Imaging. Computational Challenges and Clinical Opportunities: Second International Workshop, Held in Conjunction with MICCAI 2010, Beijing, China, September 20, 2010, Revised Selected Papers, volume 6668 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-25718-6. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/content/
                                        978-3-642-25718-6.

Zaman:2011:VISa

[ZRP^{+}11a] Halimah Badioze Zaman, Peter Robinson, Maria Petrou, Patrick Olivier, Timothy K. Shih, and et al, editors. Visual Informatics: Sustaining Research and Innovations: Second International Visual Informatics Conference, IVIC 2011, Selangor, Malaysia, November 9–11, Proceedings, Part I, volume 7066 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25190-0. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25190-0.

Zaman:2011:VISb

[ZRP^{+}11b] Halimah Badioze Zaman, Peter Robinson, Maria Petrou, Patrick Olivier, Timothy K. Shih, and et al, editors. Visual Informatics: Sustaining Research and Innovations: Second International Visual Informatics Conference, IVIC 2011, Selangor, Malaysia, November 9–11, Proceedings, Part II, volume 7067 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CODEN LNCSD9. ISBN 3-642-25199-4. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic).  URL http://www.
springerlink.com/content/
978-3-642-25199-3.